OpenStudy (anonymous):

A vacant city lot is being turned into a neighborhood gardden. The neighbors want to fence a triangular section of the lot and plant fowers there. The longest side of the trianglular section is 7 feet shorter than twice the shortest side. The thrird side is 6 feet longer than the shortest side. The permeter is 59 feet. How long is each side?

OpenStudy (anonymous):

Help?

myininaya (myininaya):

how can the longest side be shorter than another side?

myininaya (myininaya):

oh wait nvm

myininaya (myininaya):

it said twice the shorter k

myininaya (myininaya):

let's call the shortest side x so the longest side=2x-7

myininaya (myininaya):

third side=6+x

myininaya (myininaya):

do you understand so far?

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok the perimeter is 59 feet so add all the sides=59 x+(2x-7)+(6+x)=59 solve this for x

myininaya (myininaya):

let me know what you get

OpenStudy (anonymous):

4x+7=59?

OpenStudy (anonymous):

i mean 4x+13=59

myininaya (myininaya):

how did you get 13?

OpenStudy (anonymous):

x+2x+7+6+x=59 4x+13=59

myininaya (myininaya):

why is 7 positive?

OpenStudy (anonymous):

4x-1=59

myininaya (myininaya):

ok gj so far

myininaya (myininaya):

so how do you solve for x?

OpenStudy (anonymous):

x=15

myininaya (myininaya):

:) now it says to find all side so what is the longest side= if x=15

OpenStudy (anonymous):

2x-7?

myininaya (myininaya):

yes what is that if x=15

OpenStudy (anonymous):

23

myininaya (myininaya):

ok now one more side and we are done the third side= what if x=15

OpenStudy (anonymous):

21

myininaya (myininaya):

ok and what x represent again?

OpenStudy (anonymous):

\[(x) +(2x-7) +(x+6) = 59 \] 15, 23, 21

OpenStudy (anonymous):

x represents the shortest side

myininaya (myininaya):

gj phinx

OpenStudy (anonymous):

^^

myininaya (myininaya):

lol

OpenStudy (anonymous):

thank you (: life savers